OpenStudy (anonymous):

A pizza 16 inches in diameter is cut into 8 congruent slices. Find the area of the sector of one piece of pizza. A. 25.12 in.² B. 31.40 in.² C. 50.24 in.² D. 200.96 in.²

OpenStudy (anonymous):

help :/

jimthompson5910 (jim_thompson5910):

what is the area of the entire circle?

jimthompson5910 (jim_thompson5910):

Formula: A = pi*r^2 A = area of circle r = radius pi = 3.14 approximately

OpenStudy (anonymous):

x/360 * pi r^2

OpenStudy (anonymous):

hey do u have anymore information

OpenStudy (anonymous):

hey do u have anymore information?

OpenStudy (anonymous):

its 25.13

OpenStudy (anonymous):

it's A. 25.12 reason you are dividing 360 degrees by 8 so that gives sector angle = 45 degree now use the value of Pi as 3.14 and use the equation: \[\theta \div 360 \times \pi \times r ^{2}\] where r= 16/2 = 8 and theta = 45 degree